OpenStudy (anonymous):

A health class takes a field trip to go canoeing. A canoe was paddled downstream with the current and traveled 45 miles in 5 hours. On its return trip, the canoe was going into the current and was paddled at the same rate. The return trip took 9.5 hours. The current was flowing at 2 miles per hour during both trips.

OpenStudy (anonymous):

Part A Write an equation that you could use to find the canoe’s speed, r, in miles per hour when it traveled with the current. Write another equation that you could use to find the canoe’s speed, r, in miles per hour when it traveled against the current. Part B Solve one of your equations from Part A to determine the canoe’s speed in still water in miles per hour.

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

the rate going was \(\frac{45}{5}=9\) and the rate returning was \(\frac{45}{9.5}=\frac{90}{19}\)

OpenStudy (anonymous):

that makes if the rowing speed is \(r\) then \(r+2=9\) and \(r-2=\frac{90}{19}\)

OpenStudy (anonymous):

Thank you so much! I got it!

OpenStudy (anonymous):

great! not sure if it is right though.. .i had second thoughts